# Hate Speech Detection with Generalizable Target-aware Fairness

### v1

@@ 3.4. Hate Speech Classifier
Enhancing HSD Performance via Imitation Learning. To prevent the filtered post embedding $\tilde{\textbf{s}}$ from the byproduct of performance degradation in HSD, we further enhance the training of the HSD classifier with imitation learning. Specifically, we ask $\text{MLP}_{hate}(\tilde{\textbf{s}}_{i})$ to mimic its response when given the unfiltered version of the post embedding $\textbf{s}_{i}$. This is achieved by aligning the predicted probability distributions over the two binary classes:
@@
where $D_{KL}(\cdot||\cdot)$ measures the Kullback-Leibler (KL) divergence between two distributions, and $\hat{y}_{i}^{\prime}=\text{sigmoid}(\text{MLP}_{hate}(\textbf{s}_{i}))$ is the prediction based on the unfiltered post embedding $\textbf{s}_{i}$. The imitation loss $\mathcal{L}_{imi}$ encourages the same hate speech classifier to emit similar decisions no matter whether the post embedding is filtered or not, thus further pushing the HSD classifier to decorrelate itself with the spurious target-related features in the texts.
@@ 3.5. Adversarial Optimization via Alternation
Alongside the fundamental goal of achieving accurate HSD, GetFair encompasses two additional components that aim at adversarial goals. The target discriminator is designed to identify associated targets for each given post embedding, while the filter function $f(\cdot)$ essentially tries to deceive it such that relevant targets cannot be confidently inferred from the debiased embedding $\tilde{\textbf{s}}_{i}$. As such, this naturally translates to an adversarial training paradigm where the filter function and target discriminator are mutual adversaries to each other.
@@ 4.1. Evaluation Datasets
MHS: The measuring hate speech (MHS) corpus was created and released by (Kennedy et al., 2020a). MHS combines social media posts from YouTube, Twitter, and Reddit that are manually labelled by crowdsource workers from Amazon Mechanical Turk (AMT). The AMT annotators have also marked the targets associated with each post. We also test GetFair’s generalization to unseen targets with two settings:
  • Setting 1: Seen targets – {Race, Origin, Age, Gender, Disability}; Unseen targets (test) – {eligion, Sexuality}
  • Setting 2: Seen targets – {ace, Origin, Gender, Religion, Sexuality}; Unseen targets – {Age, Disability}
@@ 4.2. Baselines and Metrics
Metrics. For evaluation, we cross-compare all methods from two perspectives: (1) the HSD effectiveness, measured by classification metrics accuracy, F1, and Area under the ROC Curve (AUC); (2) the target-specific fairness, measured by nFPED, nFNED, and HF as defined in Eq.(1) and Eq.(2).
@@
Implementation Notes. For fairness, we adopt the same pretrained encoder, BERT [^3] We adopt the base version from https://huggingface.co/transformers/v3.1.0/model_doc/bert.html. (Kenton and Toutanova, 2019) across all the methods tested for embedding the texts in each post, which is a popular and performant choice in the HSD literature (Bhandari et al., 2023; Gupta et al., 2023). For GetFair, by default we set discriminator loss coefficient $\lambda=0.9$ and the imitation loss weight $\gamma=3$ on both datasets according to the contributions and magnitudes of the loss terms. Also, we set $\mu$ to $0.9$ and $0.5$, and $K$ to $1$ and $5$ respectively for Jigsaw and MHS. $N$ and $N^{\prime}$ are set to $1$ and $5$ for optimizing GetFair, respectively. We use hidden dimension $d=256$ on both datasets. The filter depth $L$ is consistently set to $1$ for optimal efficiency as we do not notice a significant performance gain from adding more layers. As for both the target discriminator and HSD classifier, we use the same 3-layer architecture with a consistent $256$ hidden dimension in their MLPs. The implementation of GetFair is released at https://github.com/xurong-liang/GetFair.
@@ 4.3. Overall Performance (RQ1)
Effectiveness in HSD. Table 1 has shown the consistently advantageous HSD effectiveness of GetFair, which has achieved the best accuracy and F1 on Jigsaw, and has the best accuracy and second-best F1 on MHS. In general, the majority of debiased HSD methods achieve the fairness objective by making a considerable sacrifice in the HSD efficacy, witnessed by the subpar accuracy of LWBC on both Jigsaw and MHS, as well as the low F1 of several models, e.g., THSD and FEAG on both Jigsaw and MHS, as well as FairReprogram on MHS. In contrast, on top of its capability of debiasing, GetFair maintains a high level of HSD effectiveness.
Fairness in HSD. On Jigsaw, GetFair yields the lowest (hence the best) nFNED and HF scores, and is on par with the strong baseline FEAG in terms of nFPED. On MHS, GetFair has scored the second-best results for both nFNED and HF, only falling behind LWBC by a small margin. In the meantime, it is worth mentioning that the strong fairness of LWBC on MHS is achieved by giving up its utility – its 50% accuracy implies the detection effectiveness is only slightly better than a random classifier. It is also noticed that, among the two metrics nFPED and nFNED, some methods are able to obtain a relatively low score on one of them, but tend to get a substantially higher score on the other. Examples include THSD and FEAG on Jigsaw, where the nFNED scores are one magnitude higher than their nFPED scores, translating into a larger number of real hateful posts being missed out by the trained classifier. Meanwhile, GetFair keeps both nFPED and nFNED low, showcasing a balanced performance. More importantly, considering that the test set contains two unseen targets, this reflects the superior generalizability of GetFair when being applied in real-world settings.
@@ 4.4. Ablation Study (RQ2)
We hereby answer RQ2 via ablation study, where we build variants of GetFair by removing/modifying one key component at a time, and the new results from the variants are recorded in Table 2. We use F1 and HF for performance demonstration, and tested with Setting 1 on both datasets. Specifically, we are interested in the contributions of imitation learning, semantic gap alignment, and the design of multi-target filters to both the HSD effectiveness and fairness. In what follows, we introduce the corresponding variants and analyze their performance implications.
@@
Figure 4. Analysis of the impact from key hyperparameters.
@@
Replacing Parameter Ensemble with Combinatorial Embedding Filters. As we put forward a filter design explicitly for the co-existence of multiple targets within one post, we compare its performance with a straightforward counterpart, i.e., simply performing sum pooling on each individually filtered post embeddings as described in Section 3.1. That is to say, Eq.(5) is updated to $\tilde{\textbf{s}}_{i}=\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}\tilde{\textbf{s}}_{i,t}$. With this change, a dramatic drop in both the HSD accuracy and fairness is observed. This has verified that, compared with combining outputs from all single-target filters, our parameter ensemble design in GetFair is more capable of preventing the filtered post embedding from the noises.
@@ 4.5. Hyperparameter Sensitivity (RQ3)
In this section, we examine the impact of some core hyperparameters, namely the rank $K$ (Section 3.1), and the discriminator weight $\lambda$, imitation weight $\gamma$, and regularization weight $\mu$ used for the synergic loss in Algorithm 1. This is done by varying the value of one hyperparameter at a time and recording the new results achieved, while all other hyperparameters are kept to the default setting. This part of the experiments are also conducted with Setting 1 on both datasets, and similar trends can be observed with Setting 2.
Impact of $\lambda$. The coefficient $\lambda$ is tuned in $\{0.1,0.3,0.5,0.7,0.9\}$. As it essentially links to the optimization of target-specific filtering, $lambda$ has a stronger impact on HF than on F1, where HF benefits from a larger value of $\lambda$. When $\lambda$ increases from $0.1$ to $0.5$, there is a rapid improvement in the classification fairness as per the HF score. When $\lambda$ is sufficiently large (i.e., $\lambda\geq 0.7$ in our case), the fairness gain remains positive but appears to be at a lower rate.
@@
Impact of $K$. The low-rank parameterization of the filter parameters generated by the hypernetwork intends to lower the memory cost while ensuring an adequate level of efficacy. Intuitively, a larger $K$ is able to preserve more expressiveness for the generated filters. On Jigsaw, a larger $K$ generally contributes to lower F1 but better HF, while the trend on MHS is on the opposite side. A possible reason is that, on Jigsaw, a stronger filter function can help remove more target-related information from the post embedding, but may also block useful features for HSD classification; while on MHS, the spurious target-related features are more implicit due to the different language styles, hence a more capable filter function can effectively filter out those noises from the post embedding to achieve a higher F1 score but also an increased false positive/negative predictions (i.e., a higher HF).
@@ 4.6. Compatibility with Other Encoders (RQ4)
GetFair is designed to be compatible with different pretrained text encoders $g(\cdot)$ as backbones. To verify this compatibility, we have tested GetFair by using two other popular pretrained language models, namely DistilGPT2 (HuggingFace, 2019) and RoBERTa (base version) (Liu et al., 2019) as $g(\cdot)$. Similar to RQ2 and RQ3, we test on both datasets under Setting 1 and report F1 and HF metrics.
From the results in Table 3, the first conclusion drawn is that GetFair is able to maintain its performance in HSD tasks when it is paired with different pretrained encoders, especially the classification fairness measured by HF. Secondly, as per F1, RoBERTa yields the highest HSD effectiveness among the three backbones, while DistilGPT2 shows limited effectiveness gain compared with the default BERT used in GetFair. We hypothesize that this is attributed to different PLMs’ model capacity. As per the PLMs tested, DistilGPT2 has a lower capacity than the BERT-base we have used (89 million vs. 110 million parameters), hence producing a similar or even lower accuracy than BERT. In the meantime, the better accuracy of RoBERTa aligns with its higher model size (125 million parameters in the base version) and capacity.
@@ 5.2. Fairness in Hate Speech Detection
A series of research has uncovered that methods in natural language processing (NLP) tasks are subject to a variety of fairness issues (Madanagopal and Caverlee, 2023; Omrani et al., 2023; Ramponi and Tonelli, 2022). As a typical application of NLP techniques, HSD is no exception for biases. Technically, in the context of HSD, the biases can come from either the source (Badjatiya et al., 2019; Caliskan et al., 2017) (e.g., authors, annotators, and data collectors) and the target (Ramponi and Tonelli, 2022; Shah et al., 2021) of online posts. In this paper, our main focus is to demote biases towards the targets of online posts.
@@ 6. Conclusion
In this paper, to address the deficiency of existing debiasing/fairness-aware HSD methods when handling unseen targets during training, we propose GetFair, which achieves generalizable target-aware fairness in HSD by adaptively generating target-specific filters via a hypernetwork instead of training individualized ones. A suite of innovative designs including low-rank parameterization, semantic gap alignment, and the imitation learning are proposed for lowering the memory cost, regularizing the generaliazability of target-specific filters, and enhancing the HSD classifier, respectively. Through a series of experiments, we have validated the effectiveness, fairness, and generalizability of GetFair, proving it to be a viable solution to ensuring target-aware fairness in HSD. Possible extensions of GetFair in our future work include discovering time-sensitive patterns (Chen et al., 2018, 2020a) for dynamic debiasing, as well as developing lightweight variants (Yin et al., 2024) of GetFair to further enhance scalability.
@@ Appendix A Dataset Statistics
Table 4 lists the key statistics of the Jigsaw dataset after being processed for the two different unseen target settings:
@@ 2 Train 11,734 76,844 Male, Jewish
Analogously, Table 5 below corresponds to the MHS dataset, also with two different unseen target settings:

### v2

@@ 3.4. Hate Speech Classifier
Enhancing HSD Performance via Imitation Learning. To prevent the filtered post embedding $\tilde{\textbf{s}}$ from the byproduct of performance degradation, we further enhance the training of the HSD classifier with imitation learning. Specifically, we ask $\text{MLP}_{hate}(\tilde{\textbf{s}}_{i})$ to mimic its response when given the unfiltered version of the post embedding $\textbf{s}_{i}$. This is achieved by aligning the predicted probability distributions over the two binary classes:
@@
where $D_{KL}(\cdot||\cdot)$ measures the Kullback-Leibler (KL) divergence between two distributions, and $\hat{y}_{i}^{\prime}=\text{sigmoid}(\text{MLP}_{hate}(\textbf{s}_{i}))$ is the prediction based on the unfiltered post embedding $\textbf{s}_{i}$. The imitation loss $\mathcal{L}_{imi}$ encourages the same hate speech classifier to emit similar decisions no matter whether the post embedding is filtered or not, thus further decorrelating the HSD classifier with the spurious target-related features in the texts.
@@ 3.5. Adversarial Optimization via Alternation
Alongside the fundamental goal of achieving accurate HSD, GetFair encompasses two additional components that have adversarial goals. The target discriminator is designed to identify associated targets for each given post embedding, while the filter function $f(\cdot)$ essentially tries to deceive it such that relevant targets cannot be confidently inferred from the debiased embedding $\tilde{\textbf{s}}_{i}$. As such, this naturally translates to an adversarial training paradigm where the filter function and target discriminator are mutual adversaries to each other.
@@ 4.1. Evaluation Datasets
MHS. The measuring hate speech (MHS) corpus was created and released by (Kennedy et al., 2020a). MHS combines social media posts from YouTube, Twitter, and Reddit that are manually labelled by crowdsource workers from Amazon Mechanical Turk (AMT). The AMT annotators have also marked the targets associated with each post. We also test GetFair’s generalization to unseen targets with two settings:
  • Setting 1: Seen targets – {Race, Origin, Age, Gender, Disability}; Unseen targets – {Religion, Sexuality}
  • Setting 2: Seen targets – {Race, Origin, Gender, Religion, Sexuality}; Unseen targets – {Age, Disability}
@@ 4.2. Baselines and Metrics
Metrics. For evaluation, we cross-compare all methods from two perspectives: (1) the HSD effectiveness, measured by classification metrics accuracy, F1, and Area under the ROC Curve (AUC); (2) the target-specific fairness, measured by nFPED, nFNED, and HF as in Eq.(1) and Eq.(2).
@@
Implementation Notes. For fairness, we adopt the same pretrained encoder, BERT [^4] We adopt the base version from https://huggingface.co/transformers/v3.1.0/model_doc/bert.html. (Kenton and Toutanova, 2019) across all the methods tested for embedding the texts in each post, which is a popular and performant choice in the HSD literature (Bhandari et al., 2023; Gupta et al., 2023). For GetFair, by default we set discriminator loss coefficient $\lambda=0.9$ and the imitation loss weight $\gamma=3$ on both datasets according to the contributions and magnitudes of the loss terms. Also, we set $\mu$ to $0.9$ and $0.5$, and $K$ to $1$ and $5$ respectively for Jigsaw and MHS. $N$ and $N^{\prime}$ are set to $1$ and $5$ for optimizing GetFair, respectively. We use hidden dimension $d=256$ on both datasets. The filter depth $L$ is consistently set to $1$ for optimal efficiency as we do not notice a significant performance gain from adding more layers. As for both the target discriminator and HSD classifier, we use the same 3-layer architecture with a consistent $256$ hidden dimension in their MLPs.
@@ 4.3. Overall Performance (RQ1)
Effectiveness in HSD. Table 1 has shown the consistently advantageous HSD effectiveness of GetFair, which has achieved the best accuracy, F1, and AUC on Jigsaw, and is either the best or second-best across these three metrics on MHS. In general, the majority of debiased HSD methods achieve the fairness objective by making a considerable sacrifice in the HSD efficacy, witnessed by the subpar accuracy of LWBC on both Jigsaw and MHS, as well as the low F1 of several models, e.g., THSD and FEAG on both Jigsaw and MHS, as well as FairReprogram on MHS. In contrast, on top of its capability of debiasing, GetFair maintains a high level of utility.
Fairness in HSD. On Jigsaw, GetFair yields highly advantageous fairness scores under both settings. Such superiority can also be observed on MHS with Setting 2. On MHS with Setting 1, GetFair has scored the second-best results for both nFNED and HF, only falling behind LWBC by a small margin. In the meantime, it is worth mentioning that the strong fairness of LWBC on MHS is achieved by giving up its utility – its 50% accuracy implies the detection effectiveness is only slightly better than a random classifier. It is also noticed that, among the two metrics nFPED and nFNED, some methods are able to obtain a relatively low score on one of them, but tend to get a substantially higher score on the other. Examples include THSD and FEAG on Jigsaw, where the nFNED scores are one magnitude higher than their nFPED scores, translating into a larger number of real hateful posts being missed out by the trained classifier. Meanwhile, GetFair keeps both nFPED and nFNED low, showcasing a balanced performance. More importantly, considering that the test set contains two unseen targets, this reflects the superior generalizability of GetFair in real-world settings.
@@ 4.4. Ablation Study (RQ2)
We hereby answer RQ2 via ablation study, where we build variants of GetFair by removing/modifying one key component at a time, and the new results from the variants are recorded in Table 2. We use F1 and HF for performance demonstration, and test with Setting 1 on both datasets. Specifically, we are interested in the contributions of imitation learning, semantic gap alignment, and the design of multi-target filters to both the HSD effectiveness and fairness. In what follows, we introduce the corresponding variants and analyze their performance implications.
@@
Figure 4. Analysis of the impact from key hyperparameters, with effectiveness and fairness metrics F1 and HF, respectively.
@@
Replacing Parameter Ensemble with Combinatorial Embedding Filters. As we put forward a filter design explicitly for the co-existence of multiple targets within one post, we compare its performance with a straightforward counterpart, i.e., simply performing sum pooling on each individually filtered post embeddings as described in Section 3.1. That is to say, Eq.(5) is updated to $\tilde{\textbf{s}}_{i}=\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}\tilde{\textbf{s}}_{i,t}$. With this change, a dramatic drop in both the HSD accuracy and fairness is observed. This has verified that, compared with combining outputs from all single-target filters, our parameter ensemble design in GetFair is more capable of preventing the filtered post embedding from noises.
@@ 4.5. Hyperparameter Sensitivity (RQ3)
In this section, we examine the impact of some core hyperparameters, namely the discriminator weight $\lambda$, imitation weight $\gamma$, and regularization weight $\mu$ used by the synergic loss in Algorithm 1, as well as the rank $K$ (Section 3.1). This is done by adjusting one hyperparameter at a time and recording the new results achieved, while all other hyperparameters are kept to the default setting. This part of the experiments are also conducted with Setting 1 on both datasets, and similar trends can be observed with Setting 2.
Impact of $\lambda$. The coefficient $\lambda$ is tuned in $\{0.1,0.3,0.5,0.7,0.9\}$. As it essentially links to the optimization of target-specific filtering, $\lambda$ has a stronger impact on HF than on F1, where HF benefits from a larger value of $\lambda$. When $\lambda$ increases from $0.1$ to $0.5$, there is a rapid improvement in the classification fairness as per the HF score. When $\lambda$ is sufficiently large (i.e., $\lambda\geq 0.7$ in our case), the fairness gain remains positive but appears to be at a lower rate.
@@
Impact of $K$. The low-rank parameterization of the filter parameters generated by the hypernetwork intends to lower the memory cost while ensuring an adequate level of efficacy. Intuitively, a larger $K$ is able to preserve more expressiveness for the generated filters. On Jigsaw, a larger $K$ generally contributes to lower F1 but better HF, while the trend on MHS is on the opposite side. A possible reason is that, on Jigsaw, a stronger filter function removes more target-related information from the post embedding, but also blocks useful features for HSD classification; while on MHS, the spurious target-related features are more implicit, hence a more capable filter function can effectively filter out those noises from the post embedding to achieve a higher F1 score but also more false positive/negative predictions (i.e., a higher HF).
@@ 4.6. Compatibility with Other Encoders (RQ4)
GetFair is designed to be compatible with different pretrained text encoders $g(\cdot)$ as backbones. To verify this compatibility, we have tested GetFair by using two other popular pretrained language models (PLMs), namely DistilGPT2 (HuggingFace, 2019) and RoBERTa (base version) (Liu et al., 2019) as $g(\cdot)$. Similar to RQ2 and RQ3, we test on both datasets under Setting 1 and report F1 and HF metrics.
From the results in Table 3, the first conclusion drawn is that GetFair is able to maintain its performance in HSD tasks when it is paired with different pretrained encoders, especially the classification fairness measured by HF. Secondly, as per F1, RoBERTa yields the highest HSD effectiveness among the three backbones, while DistilGPT2 shows limited effectiveness gain compared with the default BERT used in GetFair. We hypothesize that this is attributed to different PLMs’ model capacity. As per the PLMs tested, DistilGPT2 has a lower capacity than the BERT-base we have used (89 million and 110 million parameters respectively), hence producing a similar or even lower accuracy than BERT. In the meantime, the better accuracy of RoBERTa aligns with its higher model size (125 million parameters in the base version) and capacity.
@@ 5.2. Fairness in Hate Speech Detection
A series of research has uncovered that methods in natural language processing (NLP) tasks are subject to a variety of fairness issues (Madanagopal and Caverlee, 2023; Omrani et al., 2023; Ramponi and Tonelli, 2022), and HSD is no exception as a typical NLP task. Technically, in the context of HSD, the biases can come from either the source (Badjatiya et al., 2019; Caliskan et al., 2017) (e.g., authors, annotators, and data collectors) and the target (Ramponi and Tonelli, 2022; Shah et al., 2021) of online posts. In this work, our main focus is to demote biases toward the targets of online posts.
@@ 6. Conclusion
In this paper, to address the deficiency of existing debiasing/fairness-aware HSD methods when handling unseen targets during training, we propose GetFair, which achieves generalizable target-aware fairness in HSD by adaptively generating target-specific filters via a hypernetwork instead of training individualized ones. A suite of innovative designs including low-rank parameterization, semantic gap alignment, and imitation learning are proposed for lowering the memory cost, regularizing the generalizability of target-specific filters, and enhancing the HSD classifier, respectively. Through a series of experiments, we have validated the effectiveness, fairness, and generalizability of GetFair, proving it to be a viable solution to ensuring target-aware fairness in HSD. Possible extensions of GetFair in our future work include discovering time-sensitive patterns (Chen et al., 2018, 2020a) for dynamic debiasing, as well as developing lightweight variants (Yin et al., 2024) of GetFair to further enhance scalability.
@@ Appendix A Dataset Statistics
Table 4 lists the key statistics of the Jigsaw dataset after being processed for the two different unseen target settings. Analogously, Table 5 below corresponds to the MHS dataset, also with two different unseen target settings.
Appendix B An Overview of Baselines